0 0
Add a comment Improve this question Transcribed image text
Answer #1

(a) P = 5.1*3.43 = 17.493 W

(B) P = 0.699 *( .331*9.8 - 3.43 * 0.743)*5.1 = 2.4787 W

(C) INCREASING